OpenStudy (raffle_snaffle):

Orca Research made an initial deposit into an account paying 3% compounded annually. Orca withdrew $4421 at the end of year 5 and made a deposit of $6629 at the end of year 7. By the end of year 12, Orca had $66251 in the account. What was Orca's initial deposit?

OpenStudy (raffle_snaffle):

@zepdrix

OpenStudy (raffle_snaffle):

I have a solution but it was wrong.=(

zepdrix (zepdrix):

Hmm thinking :d

zepdrix (zepdrix):

\[\large\rm A(t)=P\left(1+\frac{r}{n}\right)^{nt}\]So we're trying to figure out this P value, the amount initially put into the account. The principle amount.

OpenStudy (raffle_snaffle):

Yessss

zepdrix (zepdrix):

And they told us that our rate, is 3% \(\rm r=.03\) And that the number of compoundings yearly, is just one, (annual compounding), \(\rm n=1\)

zepdrix (zepdrix):

So our formula simplifies down a little bit,\[\large\rm A(t)=P(1.03)^t\]Ok with that part? :o

OpenStudy (raffle_snaffle):

Okay

OpenStudy (raffle_snaffle):

yes I know how you get that

OpenStudy (raffle_snaffle):

F = P(1+i)^n

OpenStudy (raffle_snaffle):

F = P(1+0.03)^n F = P(1.03)^n

zepdrix (zepdrix):

So then we have some weird stuff going on. At t=5 years we have,\[\large\rm A(5)=P(1.03)^5\]But notice that 4421 was `taken out of` the account at this point,\[\large\rm A(5)=P(1.03)^5-4421\]

zepdrix (zepdrix):

Hmm I think it gets a little tricky after that, thinking :p

OpenStudy (raffle_snaffle):

Here let me show you what I got as you are thinking.

zepdrix (zepdrix):

k

OpenStudy (raffle_snaffle):

[[P(1.03)^5-4421]*(1.03)^2+6629]*(1.03)^5 = 66251

OpenStudy (raffle_snaffle):

shoot that doesn't work...

zepdrix (zepdrix):

Hmm that's what I'm coming up with also... good good good. good job. But it's leading you to an incorrect answer?

OpenStudy (raffle_snaffle):

hold on

OpenStudy (raffle_snaffle):

P = 44890.72

OpenStudy (raffle_snaffle):

Is that what you got?

zepdrix (zepdrix):

44890.72825 ya

OpenStudy (raffle_snaffle):

It's a solution I even checked....

OpenStudy (raffle_snaffle):

I did this last time and got it wrong. We have three attempts per problem. My logic is correct right?

zepdrix (zepdrix):

Yes, but it's possible we're misinterpreting something in here :) thinking

OpenStudy (raffle_snaffle):

It's not as straightforward as trying to find the future value. When finding the present value it's important to remember that anything that happens on the time line has not happened to the present value, but it did happen to the future value. We have remember that as we take the future value and work backwards. A good way to check, once you have your present value, is calculate the future value based on the cash flow. It should be approximately equal to the future value given. If you want me to check your equation, just let me know.

OpenStudy (raffle_snaffle):

This is what my instructor said over an email after I ran her my last solution.

zepdrix (zepdrix):

are you timed? I just need a few minutes to try this other approach

OpenStudy (raffle_snaffle):

I am not timed. lol

zepdrix (zepdrix):

the teachers explanation doesn't make sense to me lol

OpenStudy (raffle_snaffle):

Haha same.

OpenStudy (raffle_snaffle):

Still thinking?

zepdrix (zepdrix):

ya

zepdrix (zepdrix):

darn that didn't work :d hmm

zepdrix (zepdrix):

ok imma go eat some food, maybe look at this later :\ sorry I'm out of ideas :c

zepdrix (zepdrix):

anyone other smarty pants online? :d

OpenStudy (raffle_snaffle):

It's okay I am going to her office hours in 30 min

zepdrix (zepdrix):

cool :)

OpenStudy (raffle_snaffle):

Thanks though